# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ Abstract
INTRODUCTION
Synthetic magnetic flux in lattice structures allows effective control of their transport properties and localization behaviors Dalibard et al. (2011). For particular values of the magnetic flux, such as $\pi$-flux, a strong localization mechanism induced by the magnetic field can be observed in two-dimensional structures, i.e., the so-called Aharonov-Bohm (AB) caging Aharonov and Bohm (1959); Vidal et al. (1998, 2000). These systems featuring flat energy bands are typically represented in terms of compact localized states Bergman et al. (2008); Flach et al. (2014); Maimaiti et al. (2019); Zhang and Jin (2020), in which the eigenstates are strictly confined to a specific region in real space due to destructive interference Rhim and Yang (2019), preventing extended transport through hopping processes. The flat bands exist in both quasi-one- and two-dimensional geometries, including the dice Zhang and Jin (2020), Creutz He et al. (2021), sawtooth Maimaiti et al. (2017), rhombic Creffield and Platero (2010); Bermudez et al. (2011), honeycomb Wu et al. (2007) and Lieb lattice Lieb (1989); Julku et al. (2016). In particular, the rhombic lattice subjected to external magnetic fields exhibits the formation of degenerate flat energy bands with localized eigenstates confined within individual unit cells, which have spurred significant theoretical and experimental advancements in condensed matter physics and quantum simulation Sugawa et al. (2018); Mukherjee et al. (2018); Di Liberto et al. (2019); Kremer et al. (2020); Longhi (2021); Cáceres-Aravena et al. (2022); Li et al. (2022); Maity et al. (2024); Nicolau et al. (2023).
@@ INTRODUCTION
RESULTS
Model and experimental realization
We consider a rhombic array of superconducting qubits with tunable nearest-neighbor (NN) couplings, as illustrated in Fig. 1A, in which a synthetic flux $\Phi=0$ or $\pi$ can be effectively generated from the configuration of positive or negative coupling strengths ±$\pm J$ by tunable couplers. The system can be described by the effective Hamiltonian for a rhombic lattice model with $L=3l+1$ sites:
@@ RESULTS
where $\sigma^{+}_{i,j}$ ($\sigma_{i,j}^{-}$) represents the raising (lowering) operator for the qubit at index $(i,j)$ with $i\in\left\{A,\uparrow,\downarrow\right\}$ shown in Fig. 1A, $\sigma_{i,j}^{+}|0\rangle^{\otimes L}=|1_{i,j}\rangle$, $\Delta_{i,j}/2\pi$ is the frequency detuning of the qubit relative to the average system frequency, and $J$ is the homogeneous amplitude of NN coupling strengths between qubits. When $\Phi=0$, achieved by configuring identical positive or negative coupling strengths and setting $\Delta_{i,j}=0$, the system in the single-excitation subspace can be mapped to an effective 1D array model with coupling strengths $\sqrt{2}J$. In this case, the Bell state $|+_{j}\rangle=\left(|1_{\uparrow,j}\rangle+|1_{\downarrow,j}\rangle\right)/\sqrt{2}$ is regarded as a single site $(+,j)$ in the 1D array by local unitary transformations Danieli et al. (2021), as depicted in Fig. 1B. This effective 1D array features extended dynamics with two dispersive energy bands and the states $|-_{j}\rangle=\left(\lvert 1_{\uparrow,j}\rangle-\lvert 1_{\downarrow,j}\rangle\right)/\sqrt{2}$ are zero-energy eigenstates as $H|-_{j}\rangle=0$, forming a flat band at zero energy, as shown in Fig. 1D. For $\Phi=\pi$, the system in the single-excitation subspace is divided into decoupled three-level (bulk) or two-level (edge) systems depicted in Fig. 1C. This configuration gives rise to an extreme localization driven by destructive interference, characterized by oscillating populations between site $(A,j)$ and its nearest-neighboring sites Li et al. (2022); Rosen et al. (2024b).
@@
Figure 2: Characteristic dynamics of the rhombic array with $\Phi=0$ and $\pi$. (A)-(B) Measured population dynamics $n_{j}$ versus normalized evolution time $Jt$ for (A) $\Phi=0$ and (B) $\Phi=\pi$. (C) Time evolution of the population for qubits located at the edge sites for $\Phi=0$ and $\pi$.
@@
Figure 3: Anti-symmetric detuning on qubit frequencies. (A) Illustration of the effective 1D model for the rhombic array with anti-symmetric detuning on qubit frequencies. In the basis of such a 1D system, the anti-symmetric detuning acts as an additional coupling, connecting the decoupled subsystems with three (two) sites in the bulk (boundary) when $\Phi=\pi$. (B)-(C) Measured population dynamics $n_{j}$ for (B) $\Phi=0$ and (C) $\Phi=\pi$, with anti-symmetric detuning $\Delta=0$, $\sqrt{2}J$ and $10J$.
@@
Here, we demonstrate the characteristic dynamics of the rhombic array consisting of $L=7$ qubits with effective fluxes $\Phi=0$ and $\Phi=\pi$, realized using tunable couplers in superconducting quantum circuits. As shown in Fig. 2, we present the evolution of measured population dynamics $n_{i,j}(t)=\langle 1_{i,j}\rvert\rho(t)\lvert 1_{i,j}\rangle$, for individual qubits as a function of the normalized evolution time $Jt$ under the two flux configurations. The coupling strength is set to $J\approx 2\pi\times 4.2~{}\mathrm{MHz}$, and the system contains $l=2$ plaquettes, corresponding to $L=7$ sites. By initially preparing the central qubit at site $(A,2)$ in its excited state $|1_{A,2}\rangle$, we examine the characteristic population dynamics of the rhombic array. For the zero-flux $\Phi=0$ case, we configure all coupling strengths to be negative by controlling the frequencies of the tunable couplers. The resulting population dynamics, shown in Fig. 2A, exhibit oscillations traversing all sites, consistent with the picture of an effective 1D array featuring spatially extended eigenstates (as depicted in Fig. 1B). For the $\pi$-flux $\Phi=\pi$ case, we reconfigure two coupling strengths in each plaquette to be positive. The dynamics in the $\pi$-flux rhombic system can be interpreted as a destructive interference across each $2\times 2$ plaquette, where any initially localized excitation will be in a superposition of a finite number of localized eigenstates and remain bounded over time Vidal et al. (1998); Martinez et al. (2023). In this regime, the degenerate eigenstates are localized at the site $(A,j)$ and adjacent sites in $\{(\uparrow,j)$, $(\uparrow,j-1)$, $(\downarrow,j)$, $(\downarrow,j-1)\}$. This localization leads to vanishing population dynamics at the edge sites $(A,1)$ and $(A,3)$ when the system is initialized in the state $|1_{A,2}\rangle$, as shown in Fig. 2B. Fig. 2C further illustrates the population dynamics at the edge site $(A,1)$ as a function of the normalized evolution time $Jt$ for both configurations, $\Phi=0$ and $\pi$.
Anti-symmetric detuning
@@
We experimentally investigate a coupler-assisted rhombic system with anti-symmetric detunings on the qubit frequencies in $l=2$ plaquettes with $L=7$ sites. A single-particle excitation is initialized at the edge site $(A,1)$, and the population dynamics of each site are tracked over normalized time $Jt$, as shown in Fig. 3 (B and C). In the absence of anti-symmetric detuning $\Delta=0$, the zero-flux rhombic system exhibits extended population dynamics across all sites, while the $\pi$-flux rhombic system demonstrates oscillatory, localized dynamics confined to the sites $\{(A,1),(\uparrow,1),(\downarrow,1)\}$. When the anti-symmetric detuning is involved, the dynamics of the zero-flux system become less extended, whereas the $\pi$-flux system exhibits increased delocalization compared to the detuning-free counterpart, which can be effectively described using a comb-like and trimer lattice model shown in Fig. 3A. For a larger detuning $\Delta=10J$, the localized behavior in the zero-flux rhombic system is further enhanced, as presented in Fig. 3B. In contrast, the $\pi$-flux system (Fig. 3C) demonstrates population dynamics that traverse the middle site $(A,j)$ of the unit cell in the trimer lattice due to the sizeable inter-cell couplings Martinez Alvarez and Coutinho-Filho (2019); Anastasiadis et al. (2022). Numerical simulations supporting these observations are provided in the Supplemental Material.
@@
We further investigate the energy spectrum and ground state of the minimal $\pi$-flux rhombic system comprising $L=4$ qubits. To unveil the energy spectrum, we perform a spectroscopy measurement on the qubit $Q_{A,1}$ by applying a driving pulse through XY control line while keeping the qubit frequencies on resonance $\Delta_{i,j}=0$ and setting the coupling strengths according to the Hamiltonian in Eq. (1). In this spectroscopy, an increase in the measured population indicates that the applied pump tone is in resonance with one of the eigenfrequencies of the rhombic system, giving a transition from $|0\rangle^{\otimes L}$ to the eigenstates in the single-excitation subspace Martinez et al. (2023); Paladugu et al. (2024). As shown in Fig. 4A, the spectroscopy reveals two population peaks corresponding to the eigenenergies $E=\sqrt{2}J$ and $-\sqrt{2}J$ for $\pi$-flux $\Phi=\pi$.
@@
We employ an adiabatic protocol to prepare the ground state of $\pi$-flux rhombic system Saxberg et al. (2022); Wang et al. (2024). The system is initialized in a simple ground state $|1_{A,1}\rangle$ where the excitation is localized at the site $(A,1)$. The frequency of this site is primarily detuned ($|\Delta_{A,1}|/J>3$) below the other qubit frequencies, and the couplings are turned off ($J=0$). Then we gradually evolve the system by ramping up the couplings to $J\approx 2\pi\times 4.2$ MHz and successively ramping down the detuning frequency to reach the final Hamiltonian of the rhombic system with $\Delta_{i,j}=0$. As shown in Fig. 4 (B and C), the rhombic system follows the instantaneous ground state during the slow adiabatic evolution, reaching the ground state with (un)equally distributed population on each site measured for $\Phi=0$ ($\Phi=\pi$), consistent with the results calculated in numerical simulation.
DISCUSSION
@@
In this work, we experimentally realize a synthetic $\pi$-flux rhombic system using tunable couplers in superconducting quantum circuits. By properly configuring the positive or negative coupling strengths between neighboring qubits by tunable couplers, we demonstrate the characteristic dynamics of a rhombic array of superconducting qubits with synthetic flux $\Phi=0$ and $\pi$. We illustrate that the dynamics of $\pi$-flux rhombic system with anti-symmetric detunings can be effectively described by a trimer lattice in specific bases, and demonstrate the system dynamics with different detunings acting as inter-cell couplings in the effective trimer lattice. Such equivalent representation deepens our understanding of flat-band physics in the $\pi$-flux rhombic system Danieli et al. (2021), and provides the possibility for the experimental platforms without tunable coupling strengths to simulate a topological trimer lattice Martinez Alvarez and Coutinho-Filho (2019); Anastasiadis et al. (2022) or quasiperiodic mosaic model Zhou et al. (2023) whose inter-cell couplings can be effectively controlled by the on-site detunings in real space. Using the adiabatic protocol, we further demonstrate the ground state preparation in such a rhombic system by ramping up the tunable couplings’ strengths and adjusting the qubit frequencies. Utilizing the synthetic $\pi$-flux system with high tunability, further efforts can be made to explore the quantum information encoding of the eigenstates in such degenerate system Röntgen et al. (2019) according to the similar protocols in recently developed dual-rail qubits Campbell et al. (2020); Teoh et al. (2023); Chou et al. (2024); Levine et al. (2024); Koottandavida et al. (2024).
@@
Acknowledgments: This work was supported by the Shenzhen Science and Technology Program (Grant No. RCBS20231211090824040), the National Natural Science Foundation of China (12204228, 12374474), the Innovation Program for Quantum Science and Technology (2021ZD0301703), the Shenzhen-Hong Kong Cooperation Zone for Technology and Innovation (HZQB-KCZYB-2020050), and Guangdong Basic and Applied Basic Research Foundation (2024A1515011714, 2022A1515110615).

### v2

@@ Abstract
Synthetic magnetic flux in lattice structures allows effective control of their transport properties and localization behaviors Dalibard et al. (2011). For particular values of the magnetic flux, such as $\pi$-flux, a strong localization mechanism induced by the magnetic field can be observed in two-dimensional structures, i.e., the so-called Aharonov-Bohm (AB) caging Aharonov and Bohm (1959); Vidal et al. (1998, 2000). These systems featuring flat energy bands Leykam et al. (2018); Danieli et al. (2024, 2021a); Kolovsky et al. (2023) are typically represented in terms of compact localized states Bergman et al. (2008); Flach et al. (2014); Maimaiti et al. (2019); Zhang and Jin (2020), in which the eigenstates are strictly confined to a specific region in real space due to destructive interference Rhim and Yang (2019), preventing extended transport through hopping processes. The flat bands exist in both quasi-one- and two-dimensional geometries, including the dice Zhang and Jin (2020), Creutz He et al. (2021), sawtooth Maimaiti et al. (2017), rhombic Creffield and Platero (2010); Bermudez et al. (2011), honeycomb Wu et al. (2007) and Lieb lattice Lieb (1989); Julku et al. (2016). In particular, the rhombic lattice subjected to external magnetic fields exhibits the formation of degenerate flat energy bands with localized eigenstates confined within individual unit cells, which have spurred significant theoretical and experimental advancements in condensed matter physics and quantum simulation Sugawa et al. (2018); Mukherjee et al. (2018); Di Liberto et al. (2019); Kremer et al. (2020); Longhi (2021); Cáceres-Aravena et al. (2022); Li et al. (2022); Maity et al. (2024); Nicolau et al. (2023).
@@
Model and experimental realization. We consider a rhombic array of superconducting qubits with tunable nearest-neighbor (NN) couplings, as illustrated in Fig. 1(a), in which a synthetic flux $\Phi=0$ or $\pi$ can be effectively generated from the configuration of positive or negative coupling strengths ±$\pm J$ by tunable couplers. The system can be described by the effective Hamiltonian for a rhombic lattice model with $L=3l+1$ sites:
@@
where $\sigma^{+}_{i,j}$ ($\sigma_{i,j}^{-}$) represents the raising (lowering) operator for the qubit at index $(i,j)$ with $i\in\left\{A,\uparrow,\downarrow\right\}$ shown in Fig. 1(a), $\sigma_{i,j}^{+}|0\rangle^{\otimes L}=|1_{i,j}\rangle$, $\Delta_{i,j}/2\pi$ is the frequency detuning of the qubit relative to the average system frequency, and $J$ is the homogeneous amplitude of NN coupling strengths between qubits. When $\Phi=0$, achieved by configuring identical positive or negative coupling strengths and setting $\Delta_{i,j}=0$, the system in the single-excitation subspace can be mapped to an effective 1D array model with coupling strengths $\sqrt{2}J$. In this case, the Bell state $|+_{j}\rangle=\left(|1_{\uparrow,j}\rangle+|1_{\downarrow,j}\rangle\right)/\sqrt{2}$ is regarded as a single site $(+,j)$ in the 1D array by local unitary transformations Danieli et al. (2021b), as depicted in Fig. 1(b). This effective 1D array features extended dynamics with two dispersive energy bands and the states $|-_{j}\rangle=\left(\lvert 1_{\uparrow,j}\rangle-\lvert 1_{\downarrow,j}\rangle\right)/\sqrt{2}$ are zero-energy eigenstates as $H|-_{j}\rangle=0$, forming a flat band at zero energy, as shown in Fig. 1(d). For $\Phi=\pi$, the system in the single-excitation subspace is divided into decoupled three-level (bulk) or two-level (edge) systems depicted in Fig. 1(c). This configuration gives rise to an extreme localization driven by destructive interference, characterized by oscillating populations between site $(A,j)$ and its nearest-neighboring sites Li et al. (2022); Rosen et al. (2024b).
@@
Figure 2: Characteristic dynamics of the rhombic array with $\Phi=0$ and $\pi$. (a)-(b) Measured population dynamics $n_{i,j}$ versus normalized evolution time $Jt$ for (a) $\Phi=0$ and (b) $\Phi=\pi$. (c) Time evolution of the population for qubits located at the edge sites for $\Phi=0$ and $\pi$.
@@
Figure 3: Anti-symmetric detuning on qubit frequencies. (a) Illustration of the effective 1D model for the rhombic array with anti-symmetric detuning on qubit frequencies. In the basis of such a 1D system, the anti-symmetric detuning acts as an additional coupling, connecting the decoupled subsystems with three (two) sites in the bulk (boundary) when $\Phi=\pi$. (b)-(c) Measured population dynamics $n_{i,j}$ for (b) $\Phi=0$ and (c) $\Phi=\pi$, with anti-symmetric detuning $\Delta=0$, $\sqrt{2}J$ and $10J$.
@@
Here, we demonstrate the characteristic dynamics of the rhombic array consisting of $L=7$ qubits with effective fluxes $\Phi=0$ and $\Phi=\pi$, realized using tunable couplers in superconducting quantum circuits. As shown in Fig. 2, we present the evolution of measured population dynamics $n_{i,j}(t)=\langle 1_{i,j}\rvert\rho(t)\lvert 1_{i,j}\rangle$, for individual qubits as a function of the normalized evolution time $Jt$ under the two flux configurations. The coupling strength is set to $J\approx 2\pi\times 4.2~{}\mathrm{MHz}$, and the system contains $l=2$ plaquettes, corresponding to $L=7$ sites. By initially preparing the central qubit at site $(A,2)$ in its excited state $|1_{A,2}\rangle$, we examine the characteristic population dynamics of the rhombic array. For the zero-flux $\Phi=0$ case, we configure all coupling strengths to be negative by controlling the frequencies of the tunable couplers. The resulting population dynamics, shown in Fig. 2(a), exhibit oscillations traversing all sites, consistent with the picture of an effective 1D array featuring spatially extended eigenstates (as depicted in Fig. 1(b)). For the $\pi$-flux $\Phi=\pi$ case, we reconfigure two coupling strengths in each plaquette to be positive. The dynamics in the $\pi$-flux rhombic system can be interpreted as a destructive interference across each $2\times 2$ plaquette, where any initially localized excitation will be in a superposition of a finite number of localized eigenstates and remain bounded over time Vidal et al. (1998); Martinez et al. (2023). In this regime, the degenerate eigenstates are localized at the site $(A,j)$ and adjacent sites in $\{(\uparrow,j)$, $(\uparrow,j-1)$, $(\downarrow,j)$, $(\downarrow,j-1)\}$. This localization leads to vanishing population dynamics at the edge sites $(A,1)$ and $(A,3)$ when the system is initialized in the state $|1_{A,2}\rangle$, as shown in Fig. 2(b). Fig. 2(c) further illustrates the population dynamics at the edge site $(A,1)$ as a function of the normalized evolution time $Jt$ for both configurations, $\Phi=0$ and $\pi$.
@@
We experimentally investigate a coupler-assisted rhombic system with anti-symmetric detunings on the qubit frequencies in $l=2$ plaquettes with $L=7$ sites. A single-particle excitation is initialized at the edge site $(A,1)$, and the population dynamics of each site are tracked over normalized time $Jt$, as shown in Fig. 3(b),(c). In the absence of anti-symmetric detuning $\Delta=0$, the zero-flux rhombic system exhibits extended population dynamics across all sites, while the $\pi$-flux rhombic system demonstrates oscillatory, localized dynamics confined to the sites $\{(A,1),(\uparrow,1),(\downarrow,1)\}$. When the anti-symmetric detuning is involved, the dynamics of the zero-flux system become less extended, whereas the $\pi$-flux system exhibits increased delocalization compared to the detuning-free counterpart, which can be effectively described using a comb-like and trimer lattice model shown in Fig. 3(a). For a larger detuning $\Delta=10J$, the localized behavior in the zero-flux rhombic system is further enhanced, as presented in Fig. 3(b). In contrast, the $\pi$-flux system (Fig. 3(c)) demonstrates population dynamics that traverse the middle site $(A,j)$ of the unit cell in the trimer lattice due to the sizeable inter-cell couplings Martinez Alvarez and Coutinho-Filho (2019); Anastasiadis et al. (2022). Numerical simulations supporting these observations are provided in the Supplemental Material.
@@
Ground state preparation. We further investigate the energy spectrum and ground state of the minimal $\pi$-flux rhombic system comprising $L=4$ qubits. To unveil the energy spectrum, we perform a spectroscopy measurement on the qubit $Q_{A,1}$ by applying a driving pulse through XY control line while keeping the qubit frequencies on resonance $\Delta_{i,j}=0$ and setting the coupling strengths according to the Hamiltonian in Eq. (1). In this spectroscopy, an increase in the measured population indicates that the applied pump tone is in resonance with one of the eigenfrequencies of the rhombic system, giving a transition from $|0\rangle^{\otimes L}$ to the eigenstates in the single-excitation subspace Martinez et al. (2023); Paladugu et al. (2024). As shown in Fig. 4(a), the spectroscopy reveals two population peaks corresponding to the eigenenergies $E=\sqrt{2}J$ and $-\sqrt{2}J$ for $\pi$-flux $\Phi=\pi$.
@@
We employ an adiabatic protocol to prepare the ground state of $\pi$-flux rhombic system Saxberg et al. (2022); Wang et al. (2024). The system is initialized in a simple ground state $|1_{A,1}\rangle$ where the excitation is localized at the site $(A,1)$. The frequency of this site is primarily detuned ($|\Delta_{A,1}|/J>3$) below the other qubit frequencies, and the couplings are turned off ($J=0$). Then we gradually evolve the system by ramping up the couplings to $J\approx 2\pi\times 4.2$ MHz and successively ramping down the detuning frequency to reach the final Hamiltonian of the rhombic system with $\Delta_{i,j}=0$. As shown in Fig. 4(b),(c), the rhombic system follows the instantaneous ground state during the slow adiabatic evolution, reaching the ground state with (un)equally distributed population on each site measured for $\Phi=0$ ($\Phi=\pi$), consistent with the results calculated in numerical simulation. The fidelity $F=\sum_{i,j}\sqrt{n_{i,j}n_{i,j}^{\mathrm{th}}}$ for the measured population distribution $n_{i,j}$ and theoretical population distribution $n^{\mathrm{th}}_{i,j}$ Shi et al. (2023); Yan et al. (2019) gives $F\approx 0.97$ ($F\approx 0.92$) for $\Phi=0$ ($\Phi=\pi$), where the dominated error is mainly from the decoherence effects during the qubits evolution (see Supplemental Material for more details and numerical simulation).
@@
Discussion. In this work, we experimentally realize a synthetic $\pi$-flux rhombic system using tunable couplers in superconducting quantum circuits. By properly configuring the positive or negative coupling strengths between neighboring qubits by tunable couplers, we demonstrate the characteristic dynamics of a rhombic array of superconducting qubits with synthetic flux $\Phi=0$ and $\pi$. We illustrate that the dynamics of $\pi$-flux rhombic system with anti-symmetric detunings can be effectively described by a trimer lattice in specific bases, and demonstrate the system dynamics with different detunings acting as inter-cell couplings in the effective trimer lattice. Such equivalent representation deepens our understanding of flat-band physics in the $\pi$-flux rhombic system Danieli et al. (2021b), and provides the possibility for the experimental platforms without tunable coupling strengths to simulate a topological trimer lattice Martinez Alvarez and Coutinho-Filho (2019); Anastasiadis et al. (2022) or quasiperiodic mosaic model Zhou et al. (2023) whose inter-cell couplings can be effectively controlled by the on-site detunings in real space. Using the adiabatic protocol, we further demonstrate the ground state preparation in such a rhombic system by ramping up the tunable couplings’ strengths and adjusting the qubit frequencies. Utilizing the synthetic $\pi$-flux system with high tunability, further efforts can be made to explore the quantum information encoding of the eigenstates in such degenerate system Röntgen et al. (2019) according to the similar protocols in recently developed dual-rail qubits Campbell et al. (2020); Teoh et al. (2023); Chou et al. (2024); Levine et al. (2024); Koottandavida et al. (2024).
@@
Acknowledgments: This work was supported by the Shenzhen Science and Technology Program (Grant No. RCBS20231211090824040), the National Natural Science Foundation of China (12204228, 12374474), the Innovation Program for Quantum Science and Technology (2021ZD0301703), the Shenzhen-Hong Kong Cooperation Zone for Technology and Innovation (HZQB-KCZYB-2020050), the Science, Technology and Innovation Commission of Shenzhen Municipality (KQTD20210811090049034), and Guangdong Basic and Applied Basic Research Foundation (2024A1515011714, 2022A1515110615).